# Typical Quantum States of the Universe are Observationally Indistinguishable

### v1

@@ 5 Philosophical Implications
We highlight some potential implications of observation typicality for philosophy of science. First, observation typicality places severe limitations to knowledge in a quantum universe. In a large universe, no experiment can distinguish the overwhelming majority of universal quantum states. This is an in-principle limitation that cannot be overcome by new technologies.
Second, the flip side is that, concerning the observable properties of the universal quantum state, we may know a lot about them because of the typicality results. For any observation, almost all universal quantum states look the same as the normalized projection $\rho_{0}$. Assuming our universe is typical among those that satisfy the PH, even though we do not know which it is exactly, we know there is uniformity in their predictions—they are very nearly the same as the probabilistic prediction of the Wentaculus density matrix. Fixing the PH nearly fixes the probabilistic predictions of typical quantum states.
It provides a new perspective on the predictive power of initial condition laws such as the PH. When the initial condition law delivers a high-dimensional subspace in the full Hilbert space, the precise choice of a typical quantum state makes almost no difference to the probabilistic prediction. For any observation, typical individual quantum states predict with nearly the same probability distribution. This raises an interesting question, which we leave to future work, whether typical choices of the initial condition, in some sense to be defined (e.g. typical subspace), lead to nearly the same probability distributions.
@@
Third, observation typicality may conflict with positivism, the idea that a statement is unscientific or meaningless if it cannot be tested experimentally, and a variable is not well-defined if it cannot be measured. Given the PBR theorem, and the various solutions to the quantum measurement problem, we have good reasons to accept that the universal quantum state is an objective feature of the universe. Due to observation typicality, however, it is undetertermined by observational evidence. This is ironic, to echo a point made by Cowan and Tumulka (2016): positivism is sometimes supported by examples from quantum mechanics, and yet quantum mechanics tells us that the central object in a quantum universe is well-defined, objective, but non-measurable, which undermines positivism.
Finally, our result has potential implications for realism about the quantum state. On realism, it is an open question what the universal quantum state represents and how best to understand its objectivity and reality. (For a survey, see Chen (2019).) Within realism, there are two main approaches. The first is to regard it as a physical thing, like a physical field, that exists as part of the basic building blocks of the universe (Albert 1996, Ney 2021). The second is to regard it as a physical law, like the Hamiltonian function in classical mechanics, that tells physical things (such as particles and other fields) how to move (Goldstein and Zanghì 2013).
If the quantum state is a physical thing, OI suggests an in-principle limit to how precisely it can be measured: we cannot gain substantive information about which quantum state is the actual one. Physical objects in classical mechanics do not have this character. They can be measured to arbitrary precision. Almost all quantum states of the universe will appear the same to us. This should be challenging to the empiricists who hold out hope that the basic building blocks of the universe should be in-principle measurable.
Perhaps the result is less surprising if the universal quantum state has the status of a physical law. Since it is underdetermined by observational evidence, we have to use theoretical virtues, such as simplicity and elegance, to pin it down. The theoretical flavor of how we choose the universal quantum state is similar to how we choose other physical laws, whose precise forms are constrained but still vastly underdetermined by observational evidence.
@@
Realism about the universal quantum state also leaves open whether it is necessarily a pure state represented by a wave function, or possibly an impure (“mixed”) one represented by a density matrix. Following Chen (2021, 2024b), we call the first option wave function realism (WFR) and the second density matrix realism (DMR). It is already known that WFR and DMR are empirically equivalent, given appropriate choices of the universal quantum states. Observation typicality suggests that it is even harder to distinguish the two theories. Not only are typical individual wave functions OI from each other, but they are also OI from the normalized projection, an impure density matrix. One may regard it as a potential reason to prefer DMR over WFR. If typical pure states $\Psi_{t_{0}}$ compatible with the Past Hypothesis do not lead to different predictions, why not just use the Wentaculus density matrix $\rho_{0}$ instead? After all, $\rho_{0}$ may be regarded as much simpler than any typical $\Psi_{t_{0}}$ even though they are nearly predictively equivalent.
Here we do not take a firm stance on these implications, merely exploring how they may be suggested by our results.
@@ 6 Conclusion
We have shown that the central object in quantum theory—the quantum state of the universe—is most likely hidden from observation, because typical quantum states in a high-dimensional Hilbert subspace $\mathscr{H}_{0}$ are observationally indistinguishable from the density matrix $\rho_{0}=P_{0}/d_{0}$ and thus from each other. In fact, no observation will yield any substantial information at all, even partial or probabilistic, about $\Psi_{t_{0}}$. That is in a sense the strongest known result about a limitation to knowledge in a quantum universe. Assuming the universe we inhabit is typical, our observational data alone will tell us very little about exactly which one it is that we inhabit. Nature is more secretive than we have realized.
Acknowledgment. We thank Sheldon Goldstein for extensive feedback and for suggesting the term “observation typicality.” For helpful discussions, we also thank Emily Adlam, David Albert, Fabio Anza, Patricio Avila Cardenas, Jeff Barrett, Jeremy Butterfield, Craig Callender, Branden Fitelson, Hans Halvorson, Alan Hájek, Tyler Hildebrand, Christopher Jarzynski, Joel Lebowitz, JB Manchak, Kerry McKenzie, Marcos Rigol, Daniel Rubio, Tony Short, Stefan Teufel, Isaac Wilhelm, Nicole Yunger Halpern, the UC San Diego Philosophy of Physics Group, and the participants at the 2024 Quantum Thermodynamics Workshop at the University of Maryland, College Park.

### v2

@@ 5 Philosophical Implications
We sketch four potential philosophical implications of observation typicality. Our aim here is exploratory, not prescriptive.
First, observation typicality imposes severe limitations on knowledge in a quantum universe. They are in-principle limitations that cannot be overcome, regardless of technological advancements, except in cases where the relevant provisos do not apply.
Second, the flip side is that observation typicality allows us to know a great deal about the observable properties of the universal quantum state. For any given observation, nearly all universal quantum states are indistinguishable from the normalized projection $\rho_{0}$. Hence, fixing the Past Hypothesis almost completely determines the probabilistic predictions of typical quantum states in $\mathscr{H}_{0}$.
@@
Third, observation typicality may conflict with positivism, the idea that a statement is unscientific or meaningless if it cannot be tested experimentally, and a variable is not well-defined if it cannot be measured. Given the PBR theorem (Pusey et al. 2012), and the various solutions to the quantum measurement problem, we have good reasons to accept that the universal quantum state $\Psi$ is an objective feature of a quantum universe, yet it is observationally underdetermined because of observation typicality. This is ironic, to echo a point made by Cowan and Tumulka (2016): positivism is sometimes defended using examples from quantum mechanics, yet quantum mechanics itself reveals that a central object in a quantum universe—the universal quantum state—is well-defined, objective, but fundamentally unmeasurable.
Finally, our results impact quantum state realism (see Chen (2019) for a survey). One view treats the universal quantum state as a physical entity—a field constituting one of the universe’s basic building blocks (Albert 1996, Ney 2021). Yet observation typicality implies limits to even imprecisely measuring it, challenging the empiricist expectation that fundamental constituents of the universe are in principle measurable (at least imprecisely). Alternatively, if we interpret the universal quantum state as a physical law (akin to the classical Hamiltonian; see Goldstein and Zanghì (2013)), we may naturally expect it to be constrained by observationaly evidence but remain vastly underdetermined, much like other physical laws.
@@
Quantum state realism leaves open the question of whether it must necessarily be a pure state, represented by a wave function, or if it could be a mixed state, represented by a density matrix. Following Chen (2021, 2024b), we refer to the first option as wave function realism (WFR) and the second as density matrix realism (DMR). Not only are typical individual wave functions OI from each other, but they are also OI from $\rho_{0}$, which is mixed. This may offer a defeasible reason to prefer DMR over WFR. If typical pure states $\Psi_{t_{0}}$ compatible with the Past Hypothesis yield nearly identical predictions as $\rho_{0}$, why not adopt the Wentaculus version of DMR and postulate the Wentaculus density matrix $\rho_{0}$ instead? After all, $\rho_{0}$ can be regarded as much simpler than any typical $\Psi_{t_{0}}$ while maintaining near predictive equivalence.
@@ 6 Conclusion
We have shown that a central object in quantum theory—the quantum state of the universe—is effectively hidden from observation if it is typical. Typical quantum states in a high-dimensional Hilbert subspace $\mathscr{H}_{0}$ are observationally indistinguishable from the density matrix $\rho_{0}=P_{0}/d_{0}$ and, consequently, from each other. No observation will yield substantial information about $\Psi_{t_{0}}$—not reliably and (unless the observation is extremely unlikely) not even unreliably. These results amount to the strongest epistemic limits we know for a quantum universe. If the wave function of our own universe is typical in $\mathscr{H}_{0}$, our observational data alone will reveal very little about which quantum state it is in. Nature, it seems, is far more secretive than we have realized.
Acknowledgment. We thank Sheldon Goldstein for his extensive feedback and for suggesting the term “observation typicality.” For helpful discussions, we are also grateful to Emily Adlam, David Albert, Fabio Anza, Patricio Avila Cardenas, Jeff Barrett, Jeremy Butterfield, Craig Callender, Jennifer Carr, Silvia De Bianchi, Branden Fitelson, Hans Halvorson, Alan Hájek, Tyler Hildebrand, Nick Huggett, Christopher Jarzynski, Andrew Jordan, Joel Lebowitz, Matthew Leifer, Barry Loewer, JB Manchak, Tim Maudlin, Travis McKenna, Kerry McKenzie, Marcos Rigol, Daniel Rubio, Charles Sebens, Shelly Yiran Shi, Tony Short, Stefan Teufel, Mordecai Waegell, Isaac Wilhelm, Nicole Yunger Halpern, the CalTech Philosophy of Science Group, the SoCal Quantum Foundations Hub, the Metro Area Philosophy of Science Group, the COSMOS research team, the UC San Diego Philosophy of Physics Group, the participants at the 2024 Quantum Thermodynamics Workshop at the University of Maryland, College Park, and the 2025 APS March Meeting.